### v1

# Irrationality of rapidly converging series: a problem of Erdős and Graham

Abstract.
Answering a question of Erdős and Graham, we show that the double exponential growth condition $\limsup_{n\to\infty}a_{n}^{1/\phi^{n}}=\infty$ for a monotonically increasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$, together with the bound $a_{n}a_{n+1}\geq n^{1+\tau}$, is sufficient for the series $\sum_{n=1}^{\infty}1/(a_{n}a_{n+1})$ to have an irrational sum; here $\phi$ denotes the golden ratio $(1+\sqrt{5})/2$ and $\tau>0$. We also provide a positive generalization to $\sum_{n=1}^{\infty}1/(a_{n}^{w_{0}}\cdots a_{n+d-1}^{w_{d-1}})$, and a negative result showing that some of its instances are essentially optimal. The original problem was autonomously solved by the AI agent Aletheia, powered by Gemini Deep Think, while the remaining material is largely a product of human-AI interactions.
1. Declaration of AI usage
The present paper resolves a problem posed by Paul Erdős and Ronald Graham [4], which subsequently appeared in [9, 1]. Beyond the mathematical result, this work is also notable for its methodology, which relies heavily on Artificial Intelligence in a collaborative effort with Google DeepMind. We outline the research timeline below:
  1. (1)
    Erdős and Graham posed Question 1 below in their well-known 1980 problem book [4, p. 64], expressing the belief that it should be solvable (perhaps by standard tools) and that some generalization should exist.
  2. (2)
    Aletheia, a custom math-research agent built on top of Gemini Deep Think (whose development was led by Tony Feng), found a solution to Question 1. From the original output, it was clear that some generalization was possible. Specifically, the proof used rough estimates such as $1.7<2$, $\log x<x$, etc.
  3. (3)
    The authors realized and proved that one can relax the original growth condition (2.1) and generalize the series (2.2); this generalization is now formulated as Part (1) of Theorem 2.
  4. (4)
    Gemini Deep Think discovered a further generalization (among many plausible variations of the hypotheses) and then, jointly with Aletheia, proved it. This generalization corresponds to Theorem 3, with $b_{n}=1$ and a stronger additional hypothesis (2.5).
  5. (5)
    The authors found counterexamples that tightly complement the previously mentioned result; these are discussed in Part (2) of Theorem 2.
  6. (6)
    Finally, inspired by Erdős’s paper [8], the authors were able to weaken the hypotheses further to obtain Theorem 3 below. Also, a more general negative result was found and formulated as Theorem 5.
Thus, the theorems and proofs in this paper are largely products of human-AI interactions; each of the steps above also depends on the ideas discovered in the previous steps. In particular, the AI discovered the formulation of Theorem 3, after the human authors presented Theorem 2. This work also represents an interesting use case for AI in that it originated when Aletheia searched over a large set of open problems and autonomously narrowed them down to those it could solve. However, we emphasize that the writing of the paper was entirely done by the (human) authors.
2. Introduction
The following question was formulated by Erdős and Graham in [4, p. 64] and posed again by Erdős several years later [9, p. 106]. We quote its self-contained formulation from the webpage Erdős problems [1].
Question 1.
Is it true that if $a_{1}<a_{2}<\cdots$ is a sequence of positive integers with
$\liminf_{n\to\infty}a_{n}^{1/2^{n}}>1,$ (2.1)
then
$\sum_{n=1}^{\infty}\frac{1}{a_{n}a_{n+1}}$ (2.2)
is necessarily irrational?
The present paper provides a positive answer to Question 1 and its natural generalizations, namely Theorems 2 and 3. The proposers of Question 1, Erdős and Graham, noted that hypothesis (2.1) “should be enough” to imply the conclusion, while remarking that “it is not too hard to show that if $a_{n}\to\infty$ rapidly enough” [4, p. 64]. Their latter remark can be inferred from [8, Theorem 1] to mean, for example, that $\lim_{n\to\infty}a_{n}^{1/2^{n}}=\infty$. They also asked “what the strongest theorem of this type” would be and our Theorem 2 below provides a satisfactory answer.
Series of the form
$\sum_{n=1}^{\infty}\frac{1}{a_{n}}$
for a strictly increasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$ are sometimes called Ahmes series. The topic of relating their irrationality solely to the sequence growth was systematically studied in [5, 8], with very special instances appearing much earlier; see [17] for some subsequent literature. On the other hand, series of the form
$\sum_{n=1}^{\infty}\frac{b_{n}}{a_{1}a_{2}\cdots a_{n}}$ (2.3)
for some positive integer sequences $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$ with $a_{n}\geq 2$ for every $n$ are known as Cantor series. Investigation of their irrationality dates back to the mid-$19$th century [2, 20, 6, 7], while more recent relevant literature is [21, 15, 14]. These results can solve particular cases of Question 1 simply by rewriting (2.2) as the Cantor series (2.3) with $b_{n}=a_{1}\cdots a_{n-2}$. In this way, a result by Hančl and Tijdeman [15, Theorem 3.2] confirms that (2.2) is not a rational number as soon as we make the following assumptions on divisibility and the “relative growth” among the terms $a_{n}$:
$a_{n}\nmid a_{1}\cdots a_{n-2}\text{ for every }n\in\mathbb{N},\quad\liminf_{n\to\infty}\frac{a_{1}\cdots a_{n-2}}{a_{n}}=0,$
$\text{and}\quad\lim_{n\to\infty}\frac{a_{1}\cdots a_{n-2}}{a_{n-1}a_{n}}=0.$
Our result below confirms and generalizes the irrationality of (2.2) with no assumptions on the relative growth.
Theorem 2.
Fix a positive integer $d$ and let $\psi>1$ be the unique positive solution to $\psi^{d}=\psi^{d-1}+1$.
  1. (1)
    If a monotonically increasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$ satisfies
    $\lim_{n\to\infty}a_{n}^{1/\psi^{n}}=\infty,$
    then
    $\sum_{n=1}^{\infty}\frac{1}{a_{n}a_{n+1}\cdots a_{n+d-1}}$ (2.4)
    is an irrational number.
  2. (2)
    For every $C\in(1,\infty)$ there exists a strictly increasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$ satisfying
    $\lim_{n\to\infty}a_{n}^{1/\psi^{n}}=C$
    such that the infinite sum (2.4) is a rational number.
When we specialize Theorem 2 by taking $d=2$, then $\psi$ becomes the golden ratio $\phi=(1+\sqrt{5})/2<2$. In particular, it implies that condition (2.1) is sufficient for the irrationality of (2.2), while $\liminf_{n\to\infty}a_{n}^{1/\phi^{n}}>1$ is not.
We believe that Theorem 2 is a novel result when $d\geq 2$. For $d=1$ we get $\psi=2$, so Part (1) is an easy consequence of Erdős’s theorem [8, Theorem 1], while the well-known Sylvester sequence provides an explicit example of a sequence as in Part (2); see the remarks in [17, Section 1]. For $d\geq 2$, Part (1) is a special case of an even more general Theorem 3 below, while Part (2) is a particular instance of Theorem 5.
Theorem 3.
Fix a positive integer $d$ and a tuple of non-negative integers $\mathbf{w}=(w_{0},w_{1},\dots,w_{d-1})$ such that $w_{d-1}\geq 1$. Set $W=\max\{w_{0},\dots,w_{d-1}\}$, and consider the unique positive real root $c_{\mathbf{w}}$ of the polynomial
$P_{\mathbf{w}}(x)\coloneq(x-1)\sum_{j=0}^{d-1}w_{j}x^{j}-Wx^{d-1}.$
Let $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$ be sequences of positive integers, with $\{a_{n}\}_{n=1}^{\infty}$ monotonically increasing. Assume that there exist real numbers $0<\eta<\tau$ such that
$b_{n}\leq n^{\eta},\quad a_{n}^{w_{0}}a_{n+1}^{w_{1}}\cdots a_{n+d-1}^{w_{d-1}}\geq n^{1+\tau}$
for all $n\in\mathbb{N}$. We further assume that
$\limsup_{n\to\infty}a_{n}^{1/c_{\mathbf{w}}^{n}}=\infty.$
Then the infinite sum
$S_{\mathbf{w}}(\{a_{n}\}_{n=1}^{\infty},\{b_{n}\}_{n=1}^{\infty})\coloneq\sum_{n=1}^{\infty}\frac{b_{n}}{a_{n}^{w_{0}}a_{n+1}^{w_{1}}\cdots a_{n+d-1}^{w_{d-1}}}$
is irrational.
Remark 4.
  1. (1)
    It is elementary to see that $P_{\mathbf{w}}(x)=0$ has a unique positive solution $c_{\mathbf{w}}$, which belongs to $(1,\infty)$. Namely, $P_{\mathbf{w}}$ is clearly negative on $(0,1]$, while writing
    $P_{\mathbf{w}}(x)=(x-1)Q(x)-W,\quad Q(x)\coloneq w_{d-1}x^{d-1}-\sum_{j=0}^{d-2}(W-w_{j})x^{j}$
    and applying Descartes’ rule of signs to $Q$, we easily conclude that $P_{\mathbf{w}}$ has precisely one root in $(1,\infty)$.
  2. (2)
    Observe that for every sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$ and numbers $\vartheta>\theta>1$ we have the implication
    $\liminf_{n\to\infty}a_{n}^{1/\vartheta^{n}}>1\quad\implies\quad\lim_{n\to\infty}a_{n}^{1/\theta^{n}}=\infty.$
    Hence, the (stronger than needed) condition
    $\lim_{n\to\infty}a_{n}^{1/c^{n}_{\mathbf{w}}}=\infty$ (2.5)
    is met whenever we have some $b>c_{\mathbf{w}}$ satisfying lim inf_n→∞ a_n^1/b^n¿1.
  3. (3)
    Erdős [8, Theorem 1] proved the irrationality of $\sum_{n=1}^{\infty}1/a_{n}$ under the assumptions a_n ≥n^1+τ for some $\tau>0$ and every $n\in\mathbb{N}$ and lim sup_n→∞ a_n^1/2^n=∞. This is implied by the case $d=1$ in Theorem 3. Numerous variants and generalizations of this theorem by Erdős exist (see e.g., [13, 12, 11]), but we could not find any concerning series of the form (2.2) or any of the series addressed by cases $d\geq 2$ of Theorem 3. However, one should note that the aforementioned result of Erdős trivially implies the instances of Theorem 3 when $c_{\mathbf{w}}\geq 2$, which happens if and only if $\sum_{j=0}^{d-1}2^{j}w_{j}\leq 2^{d-1}W$.
The following theorem provides a negative counterpart to Theorem 3.
Theorem 5.
Fix a positive integer $d$ and a tuple of non-negative integers $\mathbf{w}=(w_{0},w_{1},\dots,w_{d-1})$ such that $w_{d-1}\geq 1$. If $\tilde{c}_{\mathbf{w}}$ is the largest positive root of the polynomial
$\widetilde{P}_{\mathbf{w}}(x)\coloneq(x-1)\sum_{j=0}^{d-1}w_{j}x^{j}-x^{d-1},$
then, for every $C\in(1,\infty)$, there exists a strictly increasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$ satisfying
$\lim_{n\to\infty}a_{n}^{1/\tilde{c}_{\mathbf{w}}^{n}}=C$ (2.6)
and
$\sum_{n=1}^{\infty}\frac{1}{a_{n}^{w_{0}}a_{n+1}^{w_{1}}\cdots a_{n+d-1}^{w_{d-1}}}\in\mathbb{Q}.$
Remark 6.
  1. (1)
    The largest real root of $\widetilde{P}_{\mathbf{w}}$ lies in $(1,\infty)$ because of $\widetilde{P}_{\mathbf{w}}(1)<0$.
  2. (2)
    Note that Theorem 5 implies that Theorem 3 is sharp in all instances when $w_{0},\dots,w_{d-1}\in\{0,1\}$.
Example 7.
When $\mathbf{w}=(1,0,2,1)$, the unique positive root of the polynomial
$P_{\mathbf{w}}(x)=(x-1)(1+2x^{2}+x^{3})-2x^{3}$
is
$c_{\mathbf{w}}=1.914\ldots,$
while the unique positive root of the polynomial
$\widetilde{P}_{\mathbf{w}}(x)=(x-1)(1+2x^{2}+x^{3})-x^{3}$
is
$\tilde{c}_{\mathbf{w}}=1.345\ldots.$
Whenever a monotonically increasing sequence of integers $\{a_{n}\}_{n=1}^{\infty}$ satisfies $a_{n}\geq n^{1+\tau}$ for some $\tau>0$ and $\limsup_{n\to\infty}a_{n}^{1/c_{\mathbf{w}}^{n}}=\infty$, we see from Theorem 3 that the sum
$\sum_{n=1}^{\infty}\frac{1}{a_{n}a_{n+2}^{2}a_{n+3}}$
is irrational. On the other hand, Theorem 5 implies that this sum can be a rational number for a strictly increasing sequence of integers $\{a_{n}\}_{n=1}^{\infty}$ such that $\lim_{n\to\infty}a_{n}^{1/\tilde{c}_{\mathbf{w}}^{n}}\in(1,\infty)$.
3. Proof of Theorem 3
We utilize a classical irrationality criterion that dates back to at least Fourier’s proof of the irrationality of the number $e$, first published in [3]. This completely elementary lemma is often attributed to Mahler [18, 19].
Lemma 8.
Let $\{z_{n}\}_{n=1}^{\infty}$ be non-negative rational numbers such that
$S\coloneq\sum_{n=1}^{\infty}z_{n}<\infty$
and $z_{n}>0$ for infinitely many $n\in\mathbb{N}$. Let $\{D_{N}\}_{N=1}^{\infty}$ be positive integers such that the value $D_{N}\sum_{n=1}^{N}z_{n}$ is an integer for every $N\in\mathbb{N}$. If $S$ is rational, then
$\liminf_{N\to\infty}D_{N}\Bigl(S-\sum_{n=1}^{N}z_{n}\Bigr)>0.$
Proof.
Write $S=p/q$ in lowest terms. Then $D_{N}\bigl(S-\sum_{n=1}^{N}z_{n}\bigr)$ is a positive rational with denominator dividing $q$, hence it is at least $1/q$. ∎
We also use a simple trick, which Erdős attributed to Borel.
Lemma 9 (cf. [8]).
If $\{\delta_{n}\}_{n=1}^{\infty}$ and $\{\mu_{n}\}_{n=1}^{\infty}$ are positive sequences with $\sum_{n=1}^{\infty}\delta_{n}<\infty$ and $\limsup_{n\to\infty}\mu_{n}=\infty$, then the set
$\mathcal{P}\coloneq\bigl\{m\in\mathbb{N}\,:\,\mu_{m+1}>(1+\delta_{m})\max_{n\leq m}\mu_{n}\bigr\}$
is infinite.
Proof.
Assuming the contrary and denoting the largest element of $\mathcal{P}$ by $n_{0}\in\mathbb{N}$, we get that every integer $n\geq n_{0}+1$ satisfies
$\mu_{n+1}\leq\mu_{n_{0}+1}\prod_{k=n_{0}+1}^{n}(1+\delta_{k})\leq\mu_{n_{0}+1}\prod_{k=1}^{\infty}(1+\delta_{k}).$
Since $\prod_{k=1}^{\infty}(1+\delta_{k})\leq\exp(\sum_{k=1}^{\infty}\delta_{k})<\infty$, we conclude that $\{\mu_{n}\}_{n=1}^{\infty}$ is bounded, which contradicts our assumptions. ∎
The following elementary lemma appears in [8]. We prove it by splitting positive integers into dyadic blocks. Given two functions $f$ and $g$, let us write $f(x)\ll g(x)$ when there exists some constant $C>0$ satisfying $|f(x)|\leq C|g(x)|$ for all $x$ in their common domain. This is the Vinogradov notation for $f=O(g)$.
Lemma 10.
Let $e\leq x_{1}\leq x_{2}\leq\cdots$ be a monotonically increasing sequence of real numbers and assume $0\leq y_{k}\leq k^{\eta}$ for some $\eta\geq 0$ and all $k$. Let $1\leq Q<R$ be integers.
  1. (1)
    If $x_{k}\geq k^{1+\tau}$ for all $k\in[Q,R]$ with some $\tau>\eta$, then _k=Q^R ykxk1xQ(τ-η)/1+τ.
  2. (2)
    If $x_{k}\geq e^{k}$ for all $k\in[Q,R]$, then _k=Q^R ykxk(logxQ)1+ηxQ.
The implicit constants here only depend on $\eta$ and $\tau$.
Proof.
For every integer $j\geq 0$ set
$I(j)\coloneq\{k\in[Q,R]:\ 2^{j}x_{Q}\leq x_{k}\leq 2^{j+1}x_{Q}\}.$
Then,
$\sum_{k=Q}^{R}\frac{y_{k}}{x_{k}}\leq\sum_{j=0}^{\infty}\sum_{k\in I(j)}\frac{y_{k}}{x_{k}}\leq\sum_{j=0}^{\infty}\frac{\sum_{k\in I(j)}y_{k}}{2^{j}x_{Q}}.$ (3.1)
First, we prove Part (1). If $k\in I(j)$, then
$y_{k}^{1+\tau}\leq k^{(1+\tau)\eta}\leq x_{k}^{\eta}\leq\left(2^{j+1}x_{Q}\right)^{\eta}.$
Thus, we have
$I(j)\leq\max I(j)\leq(2^{j+1}x_{Q})^{1/({1+\tau})}$
and
$\sum_{k\in I(j)}y_{k}\leq(2^{j+1}x_{Q})^{\eta/({1+\tau})}\#I(j)\leq(2^{j+1}x_{Q})^{(1+\eta)/({1+\tau})}.$
From (3.1) it follows that
$\sum_{k=Q}^{R}\frac{y_{k}}{x_{k}}\leq\sum_{j=0}^{\infty}\frac{(2^{j+1}x_{Q})^{(1+\eta)/({1+\tau})}}{2^{j}x_{Q}}\ll\frac{1}{x_{Q}^{(\tau-\eta)/({1+\tau})}}\sum_{j=0}^{\infty}\biggl(\frac{1}{2^{(\tau-\eta)/({1+\tau})}}\biggr)^{j}.$
Since $\tau>\eta$, the geometric series converges, giving (1).
Now we turn to Part (2). If $k\in I(j)$, then $e^{k}\leq x_{k}\leq 2^{j+1}x_{Q}$ and hence,
$k\leq\log(2^{j+1}x_{Q})\ll j+\log x_{Q}.$
Therefore, $y_{k}\leq k^{\eta}\ll(j+\log x_{Q})^{\eta}$ and
$\sum_{k\in I(j)}y_{k}\ll(j+\log x_{Q})^{\eta}\#I(j)\ll(j+\log x_{Q})^{1+\eta}.$
Since $\sum_{j=0}^{\infty}(j+L)^{1+\eta}/2^{j}\ll L^{1+\eta}$ for $L\geq 1$, (3.1) gives
$\sum_{k=Q}^{R}\frac{y_{k}}{x_{k}}\ll\sum_{j=0}^{\infty}\frac{(j+\log x_{Q})^{1+\eta}}{2^{j}x_{Q}}\ll\frac{(\log x_{Q})^{1+\eta}}{x_{Q}}.\qed$
For the rest of this section, we take sequences $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$, and assume the hypotheses of Theorem 3. We set $c\coloneq c_{\mathbf{w}}$ and
$\mu_{n}\coloneq\frac{\log a_{n}}{c^{n}},$ (3.2)
so that we have
$\limsup_{n\to\infty}\mu_{n}=\infty.$ (3.3)
It will be convenient to shift the indices from the original theorem formulation and denote
$x_{n}$ $\coloneq a_{n-d+1}^{w_{0}}a_{n-d+2}^{w_{1}}\cdots a_{n}^{w_{d-1}}=\exp\Bigl(\sum_{j=0}^{d-1}w_{j}c^{n-d+1+j}\mu_{n-d+1+j}\Bigr),$ (3.4)
$y_{n}$ $\coloneq b_{n-d+1}$
for every $n\geq d$, so that
$S\coloneq S_{\mathbf{w}}(\{a_{n}\}_{n=1}^{\infty},\{b_{n}\}_{n=1}^{\infty})=\sum_{n=d}^{\infty}\frac{y_{n}}{x_{n}}.$
From the theorem hypotheses we have
$y_{n}\leq n^{\eta}$ (3.5)
for all $n\in\mathbb{N}$. We also know that $x_{n}\geq(n-d+1)^{1+\tau}\geq n^{1+\tau^{\prime}}$ for $\tau^{\prime}\in(\eta,\tau)$ and large $n$. By writing $\tau$ in place of $\tau^{\prime}$ we can immediately assume that
$x_{n}\geq n^{1+\tau}$ (3.6)
for all sufficiently large $n$. Setting
$\delta_{n}\coloneq\frac{1}{n^{2}},$ (3.7)
we see that the peak-index set $\mathcal{P}$ in Lemma 9 is infinite and (3.3) implies
$\lim_{\mathcal{P}\ni n\to\infty}\mu_{n+1}=\infty.$
We also let
$D_{N}\coloneq\prod_{k=1}^{N}a_{k}^{W}=\exp\Bigl(W\sum_{k=1}^{N}c^{k}\mu_{k}\Bigr)$
and consider the tail sum
$r_{N}\coloneq\sum_{n=N+1}^{\infty}\frac{y_{n}}{x_{n}}.$ (3.8)
For the purpose of proving Theorem 3, it suffices to establish the following analytic fact; we note that its proof will not require the assumption that $a_{i}$ or $w_{i}$ are integers.
Proposition 11.
Under the above hypotheses, we have $\liminf_{N\to\infty}D_{N}r_{N}=0$.
Proof of Theorem 3, assuming Proposition 11.
Each of the numbers $x_{d},\ldots,x_{N}$clearly divides $D_{N}$, so
$D_{N}\Bigl(\frac{y_{d}}{x_{d}}+\cdots+\frac{y_{N}}{x_{N}}\Bigr)\in\mathbb{Z}.$
Mahler’s criterion (Lemma 8) can be applied to the series with terms
$z_{n}\coloneq\begin{cases}0&\text{for }1\leq n\leq d-1,\\ y_{n}/x_{n}&\text{for }n\geq d\end{cases}$
and combined with Proposition 11 to conclude that $S$ is an irrational number. ∎
Let us make the following observation regarding “local peaks”; this is one of the crucial ingredients in our adaptation of the proof from [8]. In particular, the validity of this inequality heavily depends on our choice of $c$.
Lemma 12.
Suppose we have indices $P,Q$ such that $Q\geq P+d-2$ and such that
$\mu_{Q+1}>(1+\delta_{Q})\max_{P\leq k\leq Q}\mu_{k}.$
Then we have
$\frac{(a_{P}a_{P+1}\cdots a_{Q})^{W}}{x_{Q+1}}\leq\Bigl(\frac{1}{a_{{Q+1}}}\Bigr)^{1/(Q^{2}+1)}.$
Proof.
Let us set $Q_{0}\coloneq Q-d+2\geq P$. Since $c$ is a root of $P_{\mathbf{w}}$, we have
$Wc^{d-1}=(c-1)\sum_{j=0}^{d-1}w_{j}c^{j},$
so that
$W\sum_{k=P}^{Q}c^{k}\leq W\sum_{k=1}^{Q}c^{k}\leq\frac{Wc^{Q+1}}{c-1}=c^{Q_{0}}\frac{Wc^{d-1}}{c-1}=c^{Q_{0}}\sum_{j=0}^{d-1}w_{j}c^{j}.$ (3.9)
Using (3.2), (3.4), (3.9), and (3.7) we can write
$\log\frac{(a_{P}a_{P+1}\cdots a_{Q})^{W}}{x_{Q+1}}$ $=W\sum_{k=P}^{Q}c^{k}\mu_{k}-\sum_{k=Q_{0}}^{Q+1}w_{k-Q_{0}}c^{k}\mu_{k}$
$=\sum_{k=P}^{Q_{0}-1}Wc^{k}\mu_{k}+\sum_{k=Q_{0}}^{Q}(\underbrace{W-w_{k-Q_{0}}}_{\geq 0})c^{k}\mu_{k}-w_{d-1}c^{Q+1}\mu_{Q+1}$
$\leq\frac{\mu_{Q+1}}{1+\delta_{Q}}\Bigl(\underbrace{W\sum_{k=P}^{Q}c^{k}-c^{Q_{0}}\sum_{j=0}^{d-2}w_{j}c^{j}}_{\leq w_{d-1}c^{Q+1}}\Bigr)-w_{d-1}c^{Q+1}\mu_{Q+1}$
$\leq\frac{-\delta_{Q}\mu_{Q+1}}{1+\delta_{Q}}w_{d-1}c^{Q+1}\leq-\frac{c^{Q+1}\mu_{Q+1}}{Q^{2}+1}.$
Exponentiating the last inequality gives
$\frac{(a_{P}\cdots a_{Q})^{W}}{x_{Q+1}}\leq\exp\Bigl(-\frac{c^{Q+1}\mu_{Q+1}}{Q^{2}+1}\Bigr)=\exp\Bigl(-\frac{\log a_{Q+1}}{Q^{2}+1}\Bigr)=a_{Q+1}^{-1/(Q^{2}+1)},$
as claimed. ∎
Now let us establish the proposition.
Proof of Proposition 11.
Let us set
$\alpha\coloneq\frac{\tau-\eta}{1+\tau}>0,\quad M\coloneq\Bigl\lceil\frac{2}{\alpha}\Bigr\rceil,$
and
$A\coloneq\max\Bigl\{MW+1,\,\prod_{n=1}^{\infty}(1+\delta_{n}),\,W\sup_{x\geq 1}\frac{x^{2}}{c^{x}}\Bigr\}.$
We consider three cases below. The main body of the proof is Case (B), while the other two cases require certain variations of the strategy from [8].
Case (A). There exist infinitely many indices $N$ such that $x_{N+1}\geq D_{N}^{M}$. Due to (3.5) and (3.6), Lemma 10 Part (1) can be applied to (3.8) with $Q=N+1$ and letting $R\to\infty$ to give
$r_{N}\ll\frac{1}{x_{N+1}^{\alpha}}.$
Hence, for every $N\in\mathbb{N}$ such that $x_{N+1}\geq D_{N}^{M}$,
$D_{N}r_{N}\ll\frac{D_{N}}{x_{N+1}^{\alpha}}\leq\frac{x_{N+1}^{1/M}}{x_{N+1}^{\alpha}}=x_{N+1}^{\,1/M-\alpha}\to 0$
as $N\to\infty$, since $1/M\leq\alpha/2$ and $x_{N+1}\to\infty$. This proves $\liminf_{N\to\infty}D_{N}r_{N}=0$ in Case (A).
Case (B). We have $a_{n}\geq e^{n}$ for every $n$. Take a large $N\in\mathcal{P}$. Since
$x_{n}\geq a_{n}^{w_{d-1}}\geq a_{n}\geq e^{n}$
for every $n$, we can apply Lemma 10 Part (2) with $Q=N+1$ and let $R\to\infty$ to obtain
$r_{N}\ll\frac{(\log x_{N+1})^{1+\eta}}{x_{N+1}}.$
Also, applying Lemma 12 with $P=1$ and $Q=N$ yields
$\frac{D_{N}}{x_{N+1}}=\frac{(a_{1}\cdots a_{N})^{W}}{x_{N+1}}\leq\Bigl(\frac{1}{a_{N+1}}\Bigr)^{1/(N^{2}+1)}.$
Combining the last two displays we get
$\log(D_{N}r_{N})\leq-\frac{\log a_{N+1}}{N^{2}+1}+(1+\eta)\log\log x_{N+1}+O(1).$
Since $x_{N+1}\leq a_{N+1}^{\sum_{i=0}^{d-1}w_{i}}$ and $\log a_{N+1}=c^{N+1}\mu_{N+1}$, we have
$\log\log x_{N+1}\leq\log\log a_{N+1}+O(1)\ll N+\log\mu_{N+1}.$
Hence,
$\log(D_{N}r_{N})\leq-\frac{c^{N+1}\mu_{N+1}}{N^{2}+1}+O(N)+O(\log\mu_{N+1}),$
which tends to -$\infty$ as $N\to\infty$ along $\mathcal{P}$, because $c>1$ and $\mu_{N+1}\to\infty$. Thus, $D_{N}r_{N}\to 0$ along $N\in\mathcal{P}$, proving the desired result in Case (B).
Case (C). Neither (A) nor (B) holds. This means
$x_{n+1}<D_{n}^{M}\ \,\text{for all sufficiently large }n$
and
$a_{n}<e^{n}\ \,\text{for infinitely many }n.$
We discard those finitely many indices $n$ for which the first of the two estimates fails, i.e., we only work with sufficiently large $n\in\mathbb{N}$.
Because of $a_{n+1}\leq x_{n+1}$, we have
$D_{n+1}=a_{n+1}^{W}D_{n}\leq(D_{n}^{M})^{W}D_{n}=D_{n}^{MW+1}.$
Since $A\geq MW+1$, we get by iteration
$D_{n+k}\leq D_{n}^{A^{k}}$ (3.10)
for sufficiently large $n$ and for $k\geq 1$.
Take any $R\in\mathcal{P}$ sufficiently large such that $c^{R+1}>R+1$ and $\mu_{R+1}>1$. This implies $\log a_{R+1}=c^{R+1}\mu_{R+1}>R+1$, so $a_{R+1}>e^{R+1}$. Let $P=P(R)<R$ be the largest index with $a_{P}<e^{P}$. Then
$a_{n}\geq e^{n}\ \,\text{for all }n\in[P+1,R+1].$ (3.11)
Also, by the standing assumptions,
$\lim_{\mathcal{P}\ni R\to\infty}P(R)=\infty.$
Moreover,
$D_{P}\leq a_{P}^{PW}\leq e^{P^{2}W}.$ (3.12)
Combined with (3.10), this gives
$D_{P+k}\leq D_{P}^{A^{k}}\leq e^{P^{2}WA^{k}}$
for all $k\geq 0$. Thus,
$\mu_{P+k}=\frac{\log a_{P+k}}{c^{P+k}}\leq\frac{\log D_{P+k}}{c^{P+k}}\leq\frac{P^{2}W}{c^{P}}\Bigl(\frac{A}{c}\Bigr)^{k}.$
Since $A\geq W\sup_{x\geq 1}x^{2}/c^{x}$, every large $P$ satisfies
$\mu_{P+k}\leq A^{k+1}$ (3.13)
for $k\geq 1$. In particular, $\mu_{R+1}\leq A^{R-P+2}$, so the fact that $\mu_{R+1}\to\infty$ as $R\to\infty$ along $\mathcal{P}$ implies
$\lim_{\mathcal{P}\ni R\to\infty}(R-P(R))=\infty.$
Hence, we may assume RP>dR-P>d.
Let us consider the set of “local peaks,”
$\mathcal{P}^{\ast}\coloneq\Bigl\{m\in[P+d,R]:\ \mu_{m+1}>1\ \text{ and }\mu_{m+1}>(1+\delta_{m})\max_{P\leq k\leq m}\mu_{k}\Bigr\}.$
Since $R\in\mathcal{P}$ and $\mu_{R+1}>1$, we have $R\in\mathcal{P}^{\ast}$, so $\mathcal{P}^{\ast}\neq\emptyset$. Let $Q=Q(R)$ be the smallest element of $\mathcal{P}^{\ast}$. Set temporarily $\bar{\mu}\coloneq\max_{P\leq k\leq P+d}\mu_{k}$. By (3.13), $\bar{\mu}\leq A^{d+1}$. By the minimality of $Q$ and the definition of $\mathcal{P}^{\ast}$, we inductively obtain
$\mu_{P+d+\ell}\leq\max\Bigl\{1,\bar{\mu}\prod_{j=P+d}^{P+d+\ell-1}(1+\delta_{j})\Bigr\}\leq A^{d+1}\prod_{n=1}^{\infty}(1+\delta_{n})\leq A^{d+2}$
for $1\leq\ell\leq Q-P-d$. Let us summarize our conclusions as
$\max_{P\leq k\leq Q}\mu_{k}\leq A^{d+2},\quad\mu_{Q+1}>(1+\delta_{Q})\max_{P\leq k\leq Q}\mu_{k},\quad\mu_{Q+1}>1,$ (3.14)
and
$\lim_{\mathcal{P}\ni R\to\infty}Q(R)=\infty.$
Now split
$D_{Q}r_{Q}=\underbrace{D_{Q}\sum_{k=Q+1}^{R}\frac{y_{k}}{x_{k}}}_{\eqcolon\Sigma_{1}(R)}+\underbrace{D_{Q}\sum_{k=R+1}^{\infty}\frac{y_{k}}{x_{k}}}_{\eqcolon\Sigma_{2}(R)}.$ (3.15)
Estimate for $\Sigma_{1}(R)$. Because of (3.11), Lemma 10 Part (2) applies and gives
$\Sigma_{1}(R)=D_{Q}\sum_{k=Q+1}^{R}\frac{y_{k}}{x_{k}}\ll\frac{D_{Q}(\log x_{Q+1})^{1+\eta}}{x_{Q+1}}.$
Thanks to (3.14) we can use Lemma 12 and combine it with (3.12) to conclude
$\log\frac{D_{Q}}{x_{Q+1}}=\log\frac{(a_{P+1}\cdots a_{Q})^{W}}{x_{Q+1}}+\log D_{P}\leq-\frac{c^{Q+1}\mu_{Q+1}}{Q^{2}+1}+P^{2}W.$
By (3.13) with k=Q+1Pk=Q+1-P, we have $\mu_{Q+1}\leq A^{Q-P+2}$. Hence,
$\log\log x_{Q+1}\leq\log(dW\log a_{Q+1})=\log(dWc^{Q+1}\mu_{Q+1})=O(Q).$
It follows that
$\log\Sigma_{1}(R)$ $\leq\log\frac{D_{Q}}{x_{Q+1}}+O(\log\log x_{Q+1})+O(1)$
$\leq-\frac{c^{Q+1}\mu_{Q+1}}{Q^{2}+1}+P^{2}W+O(Q)$
$\leq-\frac{c^{Q+1}}{Q^{2}+1}+O(Q^{2})$
for sufficiently large $P<Q\leq R$. The last expression tends to -$\infty$ as $\mathcal{P}\ni R\to\infty$, since then also $Q=Q(R)\to\infty$. Hence, $\Sigma_{1}(R)\to 0$.
Estimate for $\Sigma_{2}(R)$. This time we can only use (3.6), so Lemma 10 Part (1) gives
$\sum_{k=R+1}^{\infty}\frac{y_{k}}{x_{k}}\ll\frac{1}{x_{R+1}^{\alpha}}.$
Thus,
$\log\Sigma_{2}(R)\leq\log D_{Q}-\alpha\log x_{R+1}+O(1).$
Using (3.14), we may write
$\log\frac{D_{Q}}{D_{P}}=W\sum_{k=P+1}^{Q}c^{k}\mu_{k}\ll WA^{d+2}\sum_{k=P+1}^{Q}c^{k}=O(c^{R})$
and so, by (3.12),
$\log D_{Q}=\log\frac{D_{Q}}{D_{P}}+\log D_{P}\leq O(c^{R})+P^{2}W\leq O(c^{R})$
for large $P<Q\leq R$. Since $\log x_{R+1}\geq\log a_{R+1}=c^{R+1}\mu_{R+1}$, we see that
$\log\Sigma_{2}(R)\leq c^{R+1}\bigl(-\alpha\mu_{R+1}+O(1)\bigr),$
which converges to -$\infty$ as $\mathcal{P}\ni R\to\infty$, since $\mu_{R+1}\to\infty$ along $\mathcal{P}$. Hence, $\Sigma_{2}(R)\to 0$.
From the splitting (3.15) we conclude $\lim_{\mathcal{P}\ni R\to\infty}D_{Q(R)}r_{Q(R)}=0$. This completes the proof of Proposition 11, and hence, also that of Theorem 3. ∎
Remark 13.
In the above proof, we can also show
$\lim_{\mathcal{P}\ni R\to\infty}(R-Q(R))=\infty,$ (3.16)
even though this was not used. An analogous claim was made without proof in [8]. Indeed, we first note from (3.10) that
$c^{R+1}\mu_{R+1}=\log a_{R+1}\leq\log D_{R+1}\leq A^{R+1-Q}\log D_{Q}.$
By the definition of $D_{N}$ we have
$\log D_{Q}=\log D_{P+d-1}+W\sum_{k=P+d}^{Q}c^{k}\mu_{k},$
while from (3.14) we see
$\sum_{k=P+d}^{Q}c^{k}\mu_{k}\leq A^{d+2}\sum_{k=P+d}^{Q}c^{k}\leq A^{d+2}c^{Q+1}/(c-1).$
We have seen above that $\log D_{Q}\ll P^{2}+c^{Q}$ and so
$c^{R+1}\mu_{R+1}\ll A^{R+1-Q}(P^{2}+c^{Q}).$
Since $c^{-(R+1)}(P^{2}+c^{Q})$ is bounded and since $\mu_{R+1}\to\infty$, we obtain (3.16).
4. Proof of Theorem 5
The following is the main technical lemma, whose proof is non-constructive. Let us fix $d$ and $\mathbf{w}$, and briefly write $S(\{a_{n}\}_{n=1}^{\infty})$ instead of $S_{\mathbf{w}}(\{a_{n}\}_{n=1}^{\infty},\{1\}_{n=1}^{\infty})$.
Lemma 14.
Fix a positive integer $d$. Let $\{\beta_{n}\}_{n=1}^{\infty}$ and $\{\gamma_{n}\}_{n=1}^{\infty}$ be monotonically increasing sequences of positive integers such that $\beta_{n}\leq\gamma_{n}$ for every $n\in\mathbb{N}$,
$\sum_{n=1}^{\infty}\frac{1}{\beta_{n}^{w_{0}}\beta_{n+1}^{w_{1}}\cdots\beta_{n+d-1}^{w_{d-1}}}<\infty,$ (4.1)
$\lim_{n\to\infty}\frac{\beta_{n}}{\gamma_{n}}=0,$ (4.2)
and
$\lim_{n\to\infty}\Bigl(\frac{\beta_{n}}{\gamma_{n+1}}\Bigr)^{w_{0}}\cdots\Bigl(\frac{\beta_{n+d-2}}{\gamma_{n+d-1}}\Bigr)^{w_{d-2}}\Bigl(\frac{\beta_{n+d-1}}{\beta_{n+d}}\Bigr)^{w_{d-1}}\beta_{n+d-1}=\infty.$ (4.3)
Then the set of infinite sums
$\bigl\{S(\{a_{n}\}_{n=1}^{\infty})\,:\,(\forall n\in\mathbb{N})(a_{n}\in[\beta_{n},\gamma_{n}]\cap\mathbb{N})\bigr\}$ (4.4)
is equal to a finite union of non-degenerate closed bounded intervals. In particular, it contains a rational number.
Proof.
Let $J_{n}\coloneq[\beta_{n},\gamma_{n}]\cap\mathbb{N}$ for every $n\in\mathbb{N}$. We only need to show that there exists an $M\in\mathbb{N}$ such that, for all indices $N\geq M$ and all choices of $a_{n}\in J_{n}$, $1\leq n\leq N$, the closed interval
$\bigl[S(a_{1},\ldots,a_{N},\gamma_{N+1},\gamma_{N+2},\ldots),\,S(a_{1},\ldots,a_{N},\beta_{N+1},\beta_{N+2},\ldots)\bigr]$ (4.5)
is equal to the union of
$\bigl[S(a_{1},\ldots,a_{N},a_{N+1},\gamma_{N+2},\ldots),\,S(a_{1},\ldots,a_{N},a_{N+1},\beta_{N+2},\ldots)\bigr]$ (4.6)
taken over $a_{N+1}\in J_{N+1}$. Afterwards, we can conclude that (4.4) is equal to the union of intervals
$\bigl[S(a_{1},\ldots,a_{M},\gamma_{M+1},\gamma_{M+2},\ldots),\,S(a_{1},\ldots,a_{M},\beta_{M+1},\beta_{M+2},\ldots)\bigr]$ (4.7)
over the finitely many choices of $a_{n}\in J_{n}$, $1\leq n\leq M$. Indeed, for an arbitrary real number $x$ from (4.7) we use the claim to inductively construct $a_{n}\in J_{n}$, $n=M+1,M+2,\ldots$, such that $x$ belongs to (4.5) for all $N\geq M$. Since $S(\{a_{n}\}_{n=1}^{\infty})$ also lies in the intervals (4.5) and their lengths shrink to zero as $N\to\infty$ thanks to (4.1), we obtain the representation $x=S(\{a_{n}\}_{n=1}^{\infty})$, proving the lemma.
To complete the proof it remains to show that, for all sufficiently large $N$ and all $a_{n}\in J_{n}$, $1\leq n\leq N+1$, $a_{N+1}\neq\gamma_{N+1}$, the left endpoint of (4.6) does not exceed the right endpoint of (4.6) with $a_{N+1}$ replaced by $a_{N+1}+1$, i.e.,
$S(a_{1},\ldots,a_{N},a_{N+1},\gamma_{N+2},\ldots)\leq S(a_{1},\ldots,a_{N},a_{N+1}+1,\beta_{N+2},\ldots).$ (4.8)
This will guarantee that the intervals (4.6) fully cover (4.5). Observe that
$\frac{1}{\gamma_{n}^{w_{0}}\gamma_{n+1}^{w_{1}}\cdots\gamma_{n+d-1}^{w_{d-1}}}\leq\frac{1}{\beta_{n}^{w_{0}}\beta_{n+1}^{w_{1}}\cdots\beta_{n+d-1}^{w_{d-1}}}$
holds for all $n\geq N+2$ and that
$\frac{1}{a_{n}^{w_{0}}\cdots a_{N+1}^{w_{N-n+1}}\gamma_{N+2}^{w_{N-n+2}}\cdots\gamma_{n+d-1}^{w_{d-1}}}\leq\frac{1}{a_{n}^{w_{0}}\cdots(a_{N+1}+1)^{w_{N-n+1}}\beta_{N+2}^{w_{N-n+2}}\cdots\beta_{n+d-1}^{w_{d-1}}}$
holds for sufficiently large $N$ and for $N-d+4\leq n\leq N+1$, simply due to $w_{d-1}>0$ and $\lim_{n\to\infty}\beta_{n+d-1}/\gamma_{n+d-1}=0$. Thus, inequality (4.8) will follow from comparing contributions of only two terms from each of the two series, those corresponding to indices n=Nd+2n=N-d+2 and n=Nd+3n=N-d+3:
$\frac{1}{a_{N-d+2}^{w_{0}}\cdots a_{N}^{w_{d-2}}a_{N+1}^{w_{d-1}}}+\frac{1}{a_{N-d+3}^{w_{0}}\cdots a_{N}^{w_{d-3}}a_{N+1}^{w_{d-2}}\gamma_{N+2}^{w_{d-1}}}$
$\leq\frac{1}{a_{N-d+2}^{w_{0}}\cdots a_{N}^{w_{d-2}}(a_{N+1}+1)^{w_{d-1}}}+\frac{1}{a_{N-d+3}^{w_{0}}\cdots a_{N}^{w_{d-3}}(a_{N+1}+1)^{w_{d-2}}\beta_{N+2}^{w_{d-1}}}.$
Rearranging the terms, this inequality can be equivalently rewritten as
$\Bigl(\frac{a_{N-d+2}}{a_{N-d+3}}\Bigr)^{w_{0}}\cdots\Bigl(\frac{a_{N}}{a_{N+1}}\Bigr)^{w_{d-2}}\Bigl(\frac{a_{N+1}}{\beta_{N+2}}\Bigr)^{w_{d-1}}a_{N+1}$
$\geq a_{N+1}\biggl(1-\Bigl(\frac{a_{N+1}}{a_{N+1}+1}\Bigr)^{w_{d-1}}\biggr)\bigg/\biggl(\Bigl(\frac{a_{N+1}}{a_{N+1}+1}\Bigr)^{w_{d-2}}-\Bigl(\frac{\beta_{N+2}}{\gamma_{N+2}}\Bigr)^{w_{d-1}}\biggr).$
On the one hand, by (4.2) and the limit
$\lim_{x\to\infty}x\left(1-\Bigl(\frac{x}{x+1}\Bigr)^{w_{d-1}}\right)=w_{d-1},$
the right-hand side is bounded from above by a finite constant that depends only on $\mathbf{w}$. On the other hand, the left-hand side is at least
$\Bigl(\frac{\beta_{N-d+2}}{\gamma_{N-d+3}}\Bigr)^{w_{0}}\cdots\Bigl(\frac{\beta_{N}}{\gamma_{N+1}}\Bigr)^{w_{d-2}}\Bigl(\frac{\beta_{N+1}}{\beta_{N+2}}\Bigr)^{w_{d-1}}\beta_{N+1},$
which, by (4.3), converges to infinity as $N\to\infty$. Thus, the left-hand side is larger than the right-hand side for all sufficiently large $N\in\mathbb{N}$, and all allowed choices of $a_{n}$, $1\leq n\leq N+1$. ∎
Lemma 14 does not seem to appear in the literature, but somewhat similar constructions have been used in the context of irrationality of Ahmes and Cantor series, for instance in relation to [16, Theorems 1–3][21, Example 5.2][15, Algorithm 3.1], and [17, Lemma 5.1].
Proof of Theorem 5.
Denote $c\coloneq\tilde{c}_{\mathbf{w}}$, so that
$(c-1)\sum_{j=0}^{d-1}w_{j}c^{j}-c^{d-1}=0.$ (4.9)
For a given $C>1$ one can choose, say,
$\beta_{n}=\lfloor C^{c^{n}+n^{2}+1}\rfloor\quad\text{and}\quad\gamma_{n}=\lfloor C^{c^{n}+n^{2}+n}\rfloor$
for every $n\in\mathbb{N}$. Conditions (4.1) and (4.2) are clearly satisfied, while the limit in (4.3) is actually
$\lim_{n\to\infty}C^{\sum_{j=0}^{d-1}w_{j}(c^{n+j}-c^{n+j+1})+c^{n+d-1}+n^{2}+O(n)}=\lim_{n\to\infty}C^{n^{2}+O(n)}=\infty,$
where we used (4.9), i.e., our choice of $c$.
Lemma 14 gives an integer sequence $\{a_{n}\}_{n=1}^{\infty}$ satisfying $\beta_{n}\leq a_{n}\leq\gamma_{n}$ for every $n\in\mathbb{N}$ and $S(\{a_{n}\}_{n=1}^{\infty})\in\mathbb{Q}$. Such a sequence clearly grows according to (2.6). A minor detail is that it only becomes strictly increasing from some index onward; we surely have $a_{n}<a_{n+1}$ for all $n$ large enough so that $\gamma_{n}<\beta_{n+1}$. However, modifying finitely many terms of $\{a_{n}\}_{n=1}^{\infty}$ does not affect the rationality of the sum $S(\{a_{n}\}_{n=1}^{\infty})$. Thus, the double exponential growth of this sequence leaves more than enough room to redefine its first finite block as $a_{n}\coloneq n$ and end up with a strictly increasing sequence that still satisfies (2.6). ∎
Acknowledgements
The third named author is supported by Mid-Career Researcher Program (RS-2023-00278510) through the National Research Foundation funded by the government of Korea, and also by KIAS Individual Grants (MG073601 and MG084001, respectively) at Korea Institute for Advanced Study. The fourth named author is supported in part by the Croatian Science Foundation under the project HRZZ-IP-2022-10-5116 (FANAP). We thank Thomas Bloom for creating and maintaining the website Erdős problems and Terence Tao for valuable comments on an earlier version of this paper. We are thankful to the authors of [10] for building the Aletheia agent and letting the authors use it.